### v2

# Korean Aspect-Based Sentiment Analysis via Implicit-Feature Alignment with Corpus Filtering

Abstract
Investigations into Aspect-Based Sentiment Analysis (ABSA) for Korean restaurant reviews are notably lacking in the existing literature. Our research proposes an intuitive and effective framework for ABSA in low-resource languages such as Korean. It optimizes prediction labels by integrating translated benchmark and unlabeled Korean data. Using a model fine-tuned on translated data, we pseudo-labeled the actual Korean NLI set. Subsequently, we applied LaBSE and MSP-based filtering to this pseudo-NLI set as implicit feature, enhancing Aspect Category Detection and Polarity determination through additional training. Incorporating dual filtering, this model bridged dataset gaps, achieving positive results in Korean ABSA with minimal resources. Through additional data injection pipelines, our approach aims to utilize high-resource data and construct effective models within communities, whether corporate or individual, in low-resource language countries. Compared to English ABSA, our framework showed an approximately 3% difference in F1 scores and accuracy. We release the dataset and our code for Korean ABSA, at this link [^1] https://github.com/namkibeom/KPC-cF.
Korean Aspect-Based Sentiment Analysis via
Implicit-Feature Alignment with Corpus Filtering
Kibeom Nam Department of Data Science, Hongik University KorABSA Laboratory, MODULABS skarlqja68@gmail.com
1 Introduction
In low-resource downstream tasks such as Korean ABSA, constraints exist in constructing ABSA systems that are socially and industrially beneficial (e.g., obtaining accurate labels and high-quality training data, building a efficient serving model). Addressing this challenge is fundamentally crucial for the practical implementation of multilingual ABSA leveraging the advantages of language models Zhang et al. (2021); Lin et al. (2023). On the other hand, ABSA utilizing Large Language Models like ChatGPT can perform labeling through prompt tuning. However, it still has limitations compared to small-scale models in terms of classifier metrics and resource for training and inference Wang et al. (2023); Wu et al. (2023).
(a) TR w/o PL-cF       (b) TR w/ PL-cF
Figure 1: t-SNE visualization of last [CLS] embeddings extracted from KR3 test set by two different experimental BaselineXLM-R encoders. Our PL-cF encourages the encoder to produce discriminable representations of different sentiment polarities.
Figure 2: A diagram illustrating the two phase of our method: (1) Fine-tuning Kor-SemEval and generate pseudo-labeled KR3, (2) Fine-tuning KR3 using baseline model selected phase 1. We illustrated the filtering process (right) for fine-tuning KR3 data. Blue arrows (left & middle) indicate that this model is used to predict best label of review.
In this study, we derive pseudo-labels for real Korean reviews using machine-translated English ABSA data, inspired by the past research Balahur and Turchi (2012); Hoshino et al. (2024). Moreover, we employ Dual filtering on the actual Korean corpus converted to implicit NLI task Hendrycks and Gimpel (2017); Sun et al. (2019); Feng et al. (2022), thereby constructing an effective framework coined as Korean ABSA using Pseudo-Classifier with Corpus Filtering (KPC-cF) for implicit-feature alignment. Through this process, we assess the impact of our constructed classifier on the practical alignment of actual reviews. We validate that the pseudo-classifier, generated through the sentence-pair approach, outperforms the single approach in translation task. Furthermore, using the model that predicts the translated dataset most effectively as a baseline, we generate pseudo-labels for actual data and conduct real-world testing of Korean ABSA. This involves subsequent fine-tuning the filtered corpus based on language-agnostic embedding similarity for review and aspect sentence pairs, along with setting a threshold for Maximum Softmax Probability (MSP) in pseudo-labels.
The main contributions of our work are:
  • This is, to our knowledge, the first approach to generating a pseudo-classifier for automatic classification of aspect-based sentiment in the actual Korean domain.
  • For actual review-based ABSA, we propose a filtered NLI corpus as implicit feature and fine-tuning framework that enables important data selection in low-resource languages on models trained with high-resource dataset.
  • A new challenging dataset of Korean ABSA, along with a KR3 and translated benchmark correlated with cross-lingual understanding.
2 Two phase of Pseudo-Classifier
2.1 Motivation and Contribution
The goal of this research is to propose a framework for achieving the best ABSA on actual data with Korean nuances through high-resource languages. Past research by Balahur and Turchi Balahur and Turchi (2012) has shown that Machine Translation (MT) systems can obtain training data for languages other than English in general sentiment classification. However, existing research using translation data for alignment and alignment-free methods Li et al. (2020); Zhang et al. (2021) inadequately address the challenge of universal knowledge transfer for linguistic subgroups like Korean. Also, although it was a different domain at Zhou et al. Zhou et al. (2021), we found it necessary to investigate whether the concept of pseudo-labels could help bridge the gap of feature $\Phi$ between translated source data $\bm{D_{S}}$ and actual target language data $\bm{D_{T}}$. Therefore, we attempted the following two phases to assess the impact of the generated pseudo-classifier, fine-tuned using translated datasets from the ABSA benchmark and pseudo-labeled actual review data, on Korean ABSA. Fig. 2 shows the two-phase pseudo-classifiers we will employ. In the first phase, similar to the findings of Hoshino et al. (2024), the optimal baseline model for Korean reasoning is identified from the pool of models trained and assessed utilizing the translation dataset. In Phase 2, we fine-tune the baseline model $\Psi_{pre}(\Phi(D_{S});\theta_{D_{S}})$, which was effective in training on $D_{S}$, by additionally incorporating pseudo-labeled actual $D_{T}$. Employing the tuned model $\Psi_{post}(\Phi(D_{T});\theta_{{D_{S}}\rightarrow D_{T}})$, we conduct predictions and evaluations on manually labeled actual Korean reviews. Throughout this process, LaBSE and confidence score filtering are performed to enhance implicit features $\Phi(D_{T})$. Detailed description of our Language Adaptation for aligned ABSA task is provided in Appendix B.
Algorithm 1 Dual Filtering
1:function Sampling($Target$, $\Psi_{pre}$)
2:     for $i=1$ to Target do
3:         $x_{s},x_{a}\leftarrow$ Target sample $i$   $\triangleright\,attach\,aspect$
4:         $score_{L}\leftarrow$ LaBSE($x_{s},x_{a}$)
5:         Add $score_{L}$ to $temp$
[…]
7:         $score_{MSP}\leftarrow$ $\textit{MSP}_{\Psi_{pre}}$($y\mid x_{s},x_{a}$)
8:         if $score_{MSP}>threshold_{MSP}$ then
9:              Add $(x_{s},x_{a},y,score_{L})$ to $batch$
10:         end if
11:     end for
12:     $avg_{L}\leftarrow\frac{1}{N}\sum_{i=1}^{N}score_{L}[i]$ in $temp$
13:     $batch\leftarrow batch$ [$batch.score_{L}>avg_{L}$]
14:     return $\Phi(batch)$ for $\Psi_{post,\,joint}$ Tuning
15:end function
2.2 LaBSE based Filtering
In this approach, we aim to extract good-quality sentences-pair from the pseudo-NLI corpus. Language Agnostic BERT Sentence Embedding model Feng et al. (2022) is a multilingual embedding model that supports 109 languages, including some Korean languages. Feng et al. (2022) suggested that the dual-encoder architecture of the LaBSE model, originally designed for machine translation in source-target language data Batheja and Bhattacharyya (2022, 2023), can be applied not only to other monolingual tasks like Semantic Textual Similarity (STS) but also to data (i.e., sentence pair-set) filtering for creating high-quality training corpora in terms of meaning equivalence. Therefore, to mitigate performance degradation caused by the linguistic gap between translated ${D_{S}}$ and actual Korean ${D_{T}}$ during fine-tuning, we introduce the following filtering method that enables the identification of meaning equivalence or connotation Ghadery et al. (2019) in actual Korean sentence-pairs, even when viewed from the perspective of model trained on bilingual translation pairs. We generate the sentence embeddings for the review text and aspect of the pseudo-NLI corpora using the LaBSE model. Then, we compute the cosine similarity between the review text and aspect sentence embeddings. After that, we extract good quality NLI sentences based on a threshold value of the similarity scores.
LaBSE scoring
Let $D_{T}=\{(x_{s}^{i},x_{a}^{i})\}_{i=1}^{N}$ be a pseudo-NLI corpus with $N$ examples, where ${x_{s}^{i}}$ and ${x_{a}^{i}}$ represents $i^{th}$ review and aspect sentence respectively. We first feed all the review sentences present in the pseudo-parallel corpus as input to the LaBSE model [^2] https://huggingface.co/sentence-transformers/LaBSE, which is a Dual encoder model with BERT-based encoding modules to obtain review sentence embeddings ($S_{i}$). The sentence embeddings are extracted as the l2 normalized [CLS] token representations from the last transformer block. Then, we feed all the aspect sentences as input to the LaBSE model to obtain aspect sentence embeddings ($A_{i}$). We then compute cosine similarity $(score_{i})$ between the review and the corresponding aspect sentence embeddings.
$S_{i}=LaBSE\left(x_{s}^{i}\right)$ (1)
$A_{i}=LaBSE\left(x_{a}^{i}\right)$ (2)
$score_{i}=cosine\_similarity\left(S_{i},A_{i}\right)$ (3)
We aimed to apply the LaBSE scoring to the actual Korean data $D_{T}$, KR3, intending to facilitate flexible learning compared to the translated data $D_{S}$, Kor-SemEval (see Appendix C.1).
Dataset Count LaBSE. S Pred. Prob
(mean) (mean)
PL (BaselinemBERT) 15.23K 0.22 0.986
PL > L0.15 $\cap$ PL > Th0.5 (BaselinemBERT) 6.77K 0.24 0.988
PL (BaselineXLM-R) 15.23K 0.21 0.69
PL > L0.15 $\cap$ PL > Th0.5 (BaselineXLM-R) 6.08K 0.24 ($\uparrow$) 0.74 ($\uparrow$)
Table 1: Number of instances and average scores (LaBSE, MSP in 4 classes) for KR3 fine-tuning set.
2.3 Confidence score Filtering
Meanwhile, we need to develop a classifier $\Psi_{post}$ capable of optimal predictions on the $D_{T}$ test, which can be considered as out-of-distribution data separate from the $D_{S}$. Drawing on previous research Arora et al. (2021), we expect that language shifts ($D_{S}\leftrightarrow D_{T}$) embody both Background and Semantic shift characteristics. To ensure robust learning in both aspect detection and sentiment classification, we introduce additional thresholding on Maximum Softmax Probability (MSP; Hendrycks and Gimpel 2017) after LaBSE-based filtering on the $D_{T}$ train set. When considering an input $x=(x_{s},x_{a})\in\mathcal{X}$ and its corresponding pseudo-label $y\in\mathcal{Y}$, the score $s(x)$ for MSP is expressed as:
$s_{\textit{MSP}}(x)=\max_{k\in\mathcal{Y}}p_{\text{model}}(y=k\mid x).$ (4)
Through this, we intended a dual scoring and filtering process to ensure that our $\Psi_{post}$ does not retrain on misplaced confidence or subpar prediction outcomes for out-of-distribution data. The algorithm for calculate scores and filter with the target $D_{T}$ batch set can be found in Algorithm 1.
Model #Training Task Type Aspect Category Polarity
Storage/Docs. Precision Recall Micro-F1 4-way acc 3-way acc Binary
BaselinemBERT+PL 4.74MB 15.23K $D_{T}$ 89.71 77.60 83.21 77.96 80.15 84.01
BaselinemBERT+PL-cF 2.53MB 6.77K $D_{T}$ 87.45 76.28 81.46 78.49 80.69 84.70
BaselineXLM-R+PL 4.60MB 15.23K $D_{T}$ 90.58 79.26 84.54 83.36 85.71 89.88
BaselineXLM-R+PL-cF 2.15MB 6.08K $D_{T}$ 89.16 79.24 83.89 83.89 86.25 90.40
BaselinemBERT+TR+PL 6.28MB 30.45K $D_{S}\sqcup D_{T}$ 92.15 78.28 84.64 80.45 82.71 86.57
BaselinemBERT+TR+PL-cF 4.07MB 21.99K $D_{S}\sqcup D_{T}$ 90.15 81.20 85.42 80.88 83.15 86.96
BaselinemBERT+TR+PL 6.28MB 30.45K $D_{S}\rightarrow D_{T}$ 91.08 77.48 83.73 78.37 80.58 84.60
BaselinemBERT+TR+PL-cF 4.07MB 21.99K $D_{S}\rightarrow D_{T}$ 90.03 77.80 83.47 78.94 81.16 85.09
BaselineXLM-R+TR+PL 6.14MB 30.45K $D_{S}\sqcup D_{T}$ 91.88 83.66 87.57 83.43 85.78 89.34
KPC-cF 3.69MB 21.30K $D_{S}\sqcup D_{T}$ 92.72 83.30 87.75 83.52 85.90 89.56
BaselineXLM-R+TR+PL 6.14MB 30.45K $D_{S}\rightarrow D_{T}$ 91.50 84.23 87.69 84.21 86.57 90.35
KPC-cF 3.69MB 21.30K $D_{S}\rightarrow D_{T}$ 92.46 84.53 88.29 84.34 86.72 90.89
Table 2: KR3 test set results for Aspect Category Detection (middle) and Aspect Category Polarity (right). We reported the computational requirements as #Training. Specifically, for Baseline+PL, it refers to total number (i.e., equivalent to Kor-SemEval) and capacity of samples from KR3 Train. The performances where filtered models outperform non-filtered ones under the same training conditions are highlighted in bold & underline. KPC-cF†, is jointly fine-tuned on the Kor-SemEval & Filtered KR3 Train. Task Type is represented by Appendix B.3.
3 Experiment
3.1 Main Experiment: KR3 Test Set
Based on the results from Kor-SemEval, we observed analogous patterns between mBERT and XLM-RBase, notwithstanding their distinct properties (see Appendix D). Accordingly, We opted for the Baseline-NLI approach (i.e., BaselinemBERT, XLM-R), which demonstrated the best performance, as the base model for Phase 2.
Main Results
To investigate the effect of features $\Phi(D_{T})$ for each corpus, we conduct baseline tuning comparisons between the PL and the PL-cF (see Tab. 1, 2 for details). The variants of our tuning framework includes:
  • Baseline+PL (Pseudo-Labeled data): Fine-tuning Baseline-NLI with pseudo-KR3.
  • Baseline+PL-cF (Corpus Filtering): Fine-tuning Baseline-NLI with the data obtained by truncating instance from pseudo-KR3, where the threshold of MSP Hendrycks and Gimpel (2017) is less than 0.5 and the cosine similarity between LaBSE embeddings is less than 0.15.
  • Baseline+TR (TRanslated data)+PL: Fine-tuning Baseline-NLI (pre-tuned or jointly fine-tuned with Kor-SemEval) using pseudo-KR3.
  • KPC-cF (Baseline+TR+PL-cF): Fine-tuning Baseline-NLI (pre-tuned or jointly fine-tuned with Kor-SemEval) using PL-cF.
Results on the KR3 test set are presented in Tab. 2 and Fig. 3. We find that the KPC-cF approach achieved adequately trained results in both subtasks for the actual korean data. The model pre-tuned with Kor-SemEval achieves the best performance in Aspect Category Detection (ACD).
Figure 3: Performance of ACD and ACP on KR3 test set during adaptive fine-tuning ($D_{T}$ and $D_{S}\rightarrow D_{T}$). Left: results with the addition of other fine-tuned BaselineXLM-R. th denotes the threshold for confidence of pseudo-labeling, and L denotes the threshold for filtering of LaBSE score; Right: BaselineXLM-R tuning compared in this paper. Blue line represents our proposed model, KPC-cF.
For Aspect Category Polarity (ACP), it performs exceptionally well in the tuning of Pseudo-Labels, especially in the Binary setting. Filtered Pseudo-Labels preserve this characteristic well and amplify the performance of all metrics within ACP.
Figure 4: Top - Maximum Probability Distribution of the Fine-Tuned Model, KPC-cF (left) vs. BaselineXLM-R+TR (right), Bottom - Maximum Probability Distribution of the BaselineXLM-R+TR with LaBSE Score Distribution, All classes (left) vs. 4 classes (right).
4 Discussion
In Phase 1, XLM-R, known for its proficiency in capturing cross-lingual representations, exhibits an underfitting tendency concerning the contextual disparities in aspect vocabulary within a single task. This can be attributed to data scarcity relative to model availability for each classifier or viewed as a limitation in single task using SPM in low-resource Korean ABSA Son et al. (2023). Nevertheless, in the NLI task, it showcases potential by outperforming mBERT, guided by "aspect". Conversely, mBERT demonstrates stable results in both single and NLI tasks, exhibiting an overall accuracy increase, particularly in the NLI task (Appendix D).
Furthermore, Phase 2 reveals that the combination of the NLI approach and translation impacts the metrics of detection in aspects. Pseudo-labels in this phase contribute to enhancing the determination of sentiment, resulting in improved classifier performance. Notably, PL-cF, unlike a mere addition to translated data, play a crucial role in maintaining and enhancing accuracy and F1 score, even with fewer samples. Essentially, the filtered pseudo-NLI set alleviates the bias of pre-trained model parameters during training and enhances performance. This impact of the quality of PL further trained on TR is manifested through the embeddings generated by our model and the ablation results via thresholding (see Fig. 1, 3).
Model Joint TR+PL $\rightarrow$ $\Delta$ TR+PL-cF
Train Set Micro-F1 4-way acc
XLM-RBase I.T + LR +16.01 +6.29
mBERT I.T + LR +13.14 +5.19
TF-IDF I.T + LR +0.19 +0.19
Table 3: Performance variation of TR+PL vs. TR+PL-cF (generated by BaselineXLM-R) utilizing Logistic Regression as a shallow feature classifier for different Input Token extraction. We combined each fine-tuning set and trained all models.
Looking at Figure 4, In BaselineXLM-R+TR, the Target None class exhibits relatively high softmax probabilities, while data with polarity show relatively low probabilities, indicating significant confidence in polarity and asymmetric learning between classes. The results of KPC-cF demonstrate that additional fine-tuning with filtered data significantly increases the maximum probability of the model for Target Polarity. Simultaneously, with a reduction of over 40% in data, it prevents biased class learning, forming similar distributions among Target classes. In addition, by providing shallow features as input to a low-complexity LR model, we further confirmed the efficacy of the inherent characteristics of the filtered data through observed variations in training outcomes (see Tab. 3). Our framework suggests improving task stability and reliability by setting an optimal threshold to adjust biased classes or data with uncertain classification outcomes. Building upon Ouyang et al. Ouyang et al. (2023), we’ve shown that filtered training data is more effective for performance than simply having a large amount of information in real-world settings.
5 Conclusion
Aspect-Based Sentiment Analysis (ABSA) has been recognized as one of the most attractive subareas in text analytics and NLP. However, obtaining high-quality or ample-size labeled data has been one of the most essential issues hindering the development of ABSA. In this paper, we addressed the language gap in ABSA by building a pseudo-classifier. This involved aligning implicit features through dual filtering, further fine-tuning Korean NLI pairs with optimal pseudo-labels from a model trained on translated data. Additionally, we presented Kor-SemEval and KR3 train (pseudo labeled & filtered), testset (Gold) composed of Korean fine-grained set. We invite the community to extend Korean ABSA by providing new datasets, trained models, evaluation results, and metrics.
Acknowledgements
We thank Chanseo Nam for manually correcting and annotating the Kor-SemEval and KR3 testset used to test our model, during his undergraduate research internship. This research was supported by Brian Impact Foundation, a non-profit organization dedicated to the advancement of science and technology for all.
Appendix A Appendix
Appendix B Background and Related Work
mBERT
Multilingual BERT is a BERT trained for multilingual tasks. It was trained on monolingual Wikipedia articles in 104 different languages. It is intended to enable mBERT finetuned in one language to make predictions for another. Azhar and Khodra Azhar and Khodra (2020) and Jafarian et al. Jafarian et al. (2021) show that mbert performs effectively in a variety of multilingual Aspect-based sentiment analysis. It is also actively used as a base model in other tasks of Korean NLP Lee et al. (2021); Park et al. (2021), but is rarely confirmed in Korean ABSA tasks. Thus, our study used the pre-trained mBERT base model with 12 layers and 12 heads (i.e., 12 transfomer encoders). This model generates a 768-dimensional vector for each word. We used the 768-dimensional vector of the Extract layer to represent the comment. Like the English language subtasks, a single Dense layer was used as the classification model.
XLM-R
XLM-RoBERTa Conneau et al. (2020) is a cross-lingual model that aims to tackle the curse-of-multilingualism problem of cross-lingual models. It is inspired by RoBERTa Liu et al. (2019), trained in up to 100 languages, and outperforms mBERT in multiple cross-lingual ABSA benchmarks Zhang et al. (2021); Phan et al. (2021); Szołomicka and Kocon (2022). However, like mBERT, Korean ABSA has yet to be actively evaluated, so we used it as a base model. We use the base version (XLM-RBase) coupled with an attention head classifier, the same optimizer as mBERT.
B.1 Task description
In ABSA, Sun et al. Sun et al. (2019) set the task as equivalent to learning subtasks 3 (Aspect Category Detection) and subtask 4 (Aspect Category Polarity) of SemEval-2014 Task 4 simultaneously. A similar approach was adopted for Korean ABSA in restaurant reviews, aiming to develop a task-specific model through comparison of two PLMs, differing only in tokenization, vocabulary size, and model parameters Conneau et al. (2020). Defining a unified-serving model using multi-label-multi-class classification from a task-oriented perspective was considered impractical due to challenges in modifying pre-training set and the ongoing injection of additional data, rendering implicit mapping unattainable Sun et al. (2020); Ahmed et al. (2022); Qin and Joty (2022). Consequently, the problem has been redefined into two BERT-based classification tasks as outlined below.
Kor-SemEval Train Aspect Polarity
서비스는 평범했고 에어컨이 없어서 가격 (price) 없음 (None)
편안한 식사를 할 수 없었습니다. 일화 (anecdotes) 없음 (None)
음식 (food) 없음 (None)
(The service was mediocre and the lack of air conditioning 분위기 (ambience) 부정 (Negative)
made for a less than comfortable meal.) 서비스 (service) 중립 (Neutral)
KR3 Train Aspect Polarity
가로수길에서 조금 멀어요 점심시간에 대기 엄청납니다 가격 (price) 없음 (None)
일행 모두 있어야 들어갈 수 있어요 맛은 보통이에요 일화 (anecdotes) 부정 (Negative)
음식 (food) 없음 (None)
(It’s a little far from Garosu-gil. There’s a huge wait during lunch time. 분위기 (ambience) 부정 (Negative)
You have to have everyone in your group to get in. The taste is average.) 서비스 (service) 없음 (None)
Table 4: Samples of Kor-SemEval and KR3 train dataset.
B.2 Classification approach
Single sentence Classification
BERT for single-sentence classification tasks. For ABSA, we fine-tune the pre-trained BERT model to train $n_{a}$ (i.e., number of aspect categories) classifiers for all aspects and then summarize the results. The input representation of the BERT can explicitly represent a pair of text sentences in a sequence of tokens. In the case of a single task, only one review text is tokenized and inputted as a sequence of tokens. A given token’s input representation is constructed by summing the corresponding token, segment, and position embeddings. For classification tasks, the first word of each sequence is a unique classification embedding [CLS]. Segment embeddings in single sentence classification use one.
Sentence-pair Classification
Based on the auxiliary sentence constructed as aspect word text, we use the sentence-pair classification approach to solve ABSA. The input representation is typically the same with the single-sentence approach. The difference is that we have to add two separator tokens [SEP], the first placed between the last token of the first sentence and the first token of the second sentence. The other is placed at the end of the second sentence after its last token. This process uses both segment embeddings (In the case of XLM-RoBERTa, an additional one is placed in the first position, resulting in a total of three segment embeddings). For the training phase in the sentence-pair classification approach, we only need to train one classifier to perform both aspect categorization and sentiment classification. Add one classification layer to the Transformer output and apply the softmax activation function.
Ensemble
Meanwhile, we additionally use a voting-based ensemble, a typical ensemble method. The ensemble can confirm generalized performance based on similarity of model results in NLI task Xu et al. (2020). So, we add separate power-mean ensemble result to identify a metric that amplifies probabilities based on the Pre-trained Language Models (PLMs). we reported the ensemble results of the top-performing models trained on NLI tasks for each PLM.
B.3 Evaluation of Language Adaptation
In this paper, we focus on the langauge adaptation of ABSA. The input data includes a set of labeled sentences from a translated source data $D_{S}=\{(x_{s}^{i},x_{a}^{i},y^{i})\}_{i=1}^{N}$ and a set of unlabeled sentences from a target language data $D_{T}=\{(x_{s}^{i},x_{a}^{i})\}_{i=1}^{N}$. Our goal is to validate our proposed filtering technique across three types of tasks: (1) training only pseudo-labeled $D_{T}$, (2) joint training using both $D_{S}$ and $D_{T}$, (3) transfer learning from $D_{S}$ to $D_{T}$.
Namely, unlike Chen et al. (2023), we fixed token embeddings and fine-tuned the transformer body & head. Contrary to Gururangan et al. (2020), we trained with $D_{S}={(x_{s}^{i},x_{a}^{i},y^{i})}_{i=1}^{N}$ and then reinitialized the classifier layer for further fine-tuning. This approach examines if the universal encoder and params $\theta$ aligned with $D_{S}$ enhance implicit alignment via pseudo-labels of refined $D_{T}$. The $\Psi_{joint}(\Phi(D_{S}\sqcup D_{T});\theta_{D_{S}\sqcup D_{T}})$ model was also evaluated, despite its inefficiency in terms of memory and computation, for data validation purposes.
Appendix C Experimental Setup
C.1 Dataset for Fine-Tuning and Test
Kor-SemEval
We translate the SemEval-2014 Task 4 Pontiki et al. (2014) dataset [^3] http://alt.qcri.org/semeval2014/task4/. Moreover, it is evaluated for Korean aspect-based sentiment analysis. The training data was machine-translated (by Google Translate), and Test data was corrected manually only for fewer than 10 instances where abnormal translations occurred after machine translation. Each sentence contains a list of aspect ${x_{a}}$ with the sentiment polarity ${y}$. Ultimately, given a sentence ${x_{s}}$ in the sentence, we need to:
  • detects the mention of an aspect ${x_{a}}$;
  • determines the positive or negative sentiment polarity ${y}$ for the detected aspect.
This setting allows us to jointly evaluate Subtask 3 (Aspect Category Detection) and Subtask 4 (Aspect Category Polarity).
KR3
Unlike the domains previously used for Korean sentiment classification Lee et al. (2020); Yang (2021); Ban (2022), Korean Restaurant Review with Ratings (KR3) is a restaurant review sentiment analysis dataset constructed through actual certified map reviews. In the case of restaurant reviews, words and expressions that evaluate positive and negative are mainly included, and real users often infer what a restaurant is like by looking at its reviews. Accordingly, Jung et al.3 constructed the KR3 dataset by crawling and preprocessing user reviews and star ratings of websites that collect restaurant information and ratings. KR3 has 388,111 positive and 70,910 negative, providing a total of 459,021 data plus 182,741 unclassified data, and distributed to Hugging Face [^4] https://huggingface.co/datasets/leey4n/KR3.
We structured our training and test datasets to match the size of Kor-SemEval. Specifically, we addressed potential biases by randomly sampling indices from the original KR3, ensuring that evaluations for a specific restaurant were non-overlapping. Additionally, we maintained an even distribution of positive, negative, and neutral (ambiguous) classes, irrespective of the aspects indicated in the original KR3. This preprocessing step aimed to capture a comprehensive representation of sentiments across diverse attributes of sentences in the dataset. Subsequently, the data were configured to suit sentence pair classification. To allocate polarity labels for each aspect within the KR3 dataset, pseudo-labeling was conducted utilizing the optimal model identified during the Kor-SemEval performance evaluation. Pseudo-labels were assigned to the KR3 training data, and post pseudo-labeling, the test data underwent manual re-labeling by researchers.
Tab. 4 shows some Kor-SemEval and KR3 training data samples. In the case of KR3, the negative aspect is better reflected. Meanwhile, while Kor-SemEval gave neutrality to mediocre service, KR3 did not give neutrality to mediocre taste. While positive and negative data have been sufficiently accumulated and reflected, the tendency for a lack of neutral data can be confirmed in advance through some samples. We have organized both Kor-SemEval and KR3 data as open-source to facilitate their use in various training and evaluation scenarios.
C.2 Hyperparameter
All experiments are conducted on two pre-trained cross-lingual models. The XLM-RoBERTa-base and BERT-base Multilingual-Cased model are fine-tuned. The number of Transformer blocks is 12, the hidden layer size is 768, the number of self-attention heads is 12, and the total number of parameters for the XLM-RoBERTa-base model is 270M, and BERT base Multilingual-Cased is 110M. When fine-tuning, we keep the dropout probability at 0.1 and the initial learning rate is 2e-5.
In the Preliminary Experiment (Appendix D), we set the batch size to 3, this setting was applied to both single and NLI tasks. The max length was set to 512, and for epochs beyond 3, no significant performance improvement was observed, so the results from epoch 2 were noted. Subsequently, in Main Experiment (Sec. 3.1), we fine-tuned with a batch size of 16, following the pattern of previous experiments Karimi et al. (2021). However, no significant performance improvement was observed within 10 epochs, so we reported the results from epoch 4. Each reported metric is the average of runs with four different random seeds to mitigate the effects of random variation in the results.
Model SemEval-14
Precision Recall Micro-F1
BERT-single 92.78 89.07 90.89
BERT-pair-NLI-M 93.15 90.24 91.67
Models trained & evaluated on Kor-SemEval
mBERT-single 92.16 77.95 84.46
XLM-RBase-single 91.01 49.37 64.01
mBERT-NLI 91.10 79.90 85.14
XLM-RBase-NLI 91.37 83.71 87.37
NLI-ensemble 93.70 81.27 87.04
Table 5: Test set results for Aspect Category Detection. We use the results reported in BERT-single and BERT-pair-NLI-M Sun et al. (2019) for English dataset together with our results.
Model SemEval-14
4-way acc 3-way acc Binary
BERT-single 83.7 86.9 93.3
BERT-pair-NLI-M 85.1 88.7 94.4
Models trained & evaluated on Kor-SemEval
mBERT-single 68.20 71.84 79.52
XLM-RBase-single 62.93 66.29 75.20
mBERT-NLI 73.95 77.90 84.87
XLM-RBase-NLI 79.41 83.66 89.98
NLI-ensemble 78.24 82.43 89.65
Table 6: Test set accuracy (%) for Aspect Category Polarity. We use the results reported in BERT-single and BERT-pair-NLI-M Sun et al. (2019) for English dataset together with our results.
C.3 Metrics
The benchmarks for SemEval-2014 Task 4 are the several best performing systems in Sun et al. Sun et al. (2019), Wang et al. Wang et al. (2016) and Pontiki et al. Pontiki et al. (2014). When evaluating Kor-SemEval and KR3 test data with subtask 3 and 4, following Sun et al. Sun et al. (2019), we also use Micro-F1 and accuracy respectively.
Appendix D Preliminary Experiment
D.1 Experiment: Kor-SemEval
We conducted evaluations for each of the mBERT-single, XLM-RBase-single, mBERT-NLI, XLM-RBase-NLI, and NLI-ensemble models. We included the results from the previous SemEval14 research and Kor-SemEval to compare and evaluate the performance in Korean.
Figure 5: Left & Middle - Maximum Probability Distribution of the BaselinemBERT+TR with LaBSE Score Distribution, All classes (left) vs. 4 classes (middle), Right - Maximum Probability Distribution of the Fine-Tuned Model, BaselinemBERT+TR.
Results
Results on Kor-SemEval are presented in Tab. 5 and Tab. 6. Similar to the SemEval results, it was confirmed that tasks converted to NLI tasks tend to be better than single tasks, with mBERT achieving better results in single and XLM-RBase in NLI. The XLM-RBase-NLI model performs best, excluding precision for aspect category detection. It also works best for aspect category polarity. The NLI-ensemble model was the best in precision but performed poorly in other metrics.
Appendix E Limitations and Future Work
Acknowledging limitations is crucial before delving into future work. While mBERT exhibits somewhat inadequate transfer effects, which can be attributed to its smaller model size. Due to pre-training solely on translation data, the scores are densely populated, resulting in limited performance improvement through additional training. Nevertheless, unlike jointly with Target None, the regression line of maximum probability for Target Polarity still trends upward according to LaBSE score (see Tab 1, Figure 5), indicating that ongoing evaluation of threshold points for the two scores can be done amidst more diverse and precise patterns. Meanwhile, we conducted direct inference on individual NLI inputs through a demo [^5] https://huggingface.co/KorABSA. Upon detecting various aspects within the reviews, we observed a slight predominance towards the "None" class within the class-wise similar probability distributions per sentence in TR training, while a more concentrated prediction tendency towards the binary class was observed in PL training. Furthermore, a tendency for one class to dominate over two or more aspect classes within review sentences was identified. This appears to stem from the class imbalance in TR and contextual differences Jin et al. (2024) within all review data (i.e., TR & PL).
Dataset Count Ratio %
(T.None) (T.None/PL)
PL > L0.15 $\cap$ PL > Th0.5 (BaselinemBERT) 2891 19.02%
PL < L0.15 $\cap$ PL > Th0.5 (BaselinemBERT) 7634 50.12%
PL > L0.15 $\cap$ PL > Th0.5 (BaselineXLM-R) 37 0.243%
PL < L0.15 $\cap$ PL > Th0.5 (BaselineXLM-R) 68 0.446%
Table 7: Quality assessment results of PL filtering by two Baseline outputs; Target None ratio is 31.1% and 0.203% lower in the PL-cF threshold.
Dataset Error Ratio %
(Error/T.None)
PL > L0.15 $\cap$ PL > Th0.5 10.81%
PL < L0.15 $\cap$ PL > Th0.5 16.18%
Table 8: Qualitative assessment results of PL filtering (BaselineXLM-R) outputs; Target None errors are 5.37% lower in the PL-cF threshold.
Our analysis of the PL-cF thresholding reveals that while it reduces the number of predicted Target None classes resulting from extensive training on None class in TR, it also qualitatively explains why the error rate for Target None can be lower (see Tab. 7, 8). Contrary to experimental results that show it is reasonable to predict high-confidence Target Polarity labels when there is high surface similarity between sentences, cases where sentence similarity is less than 0.15 but predicted as Target None with confidence greater than 0.5 can hinder the model’s understanding of deeper semantic relationships. For instance, issues related to hygiene and cleanliness are not detected through the overall atmosphere, and in cases where multiple aspects overlap, the model fails to distinguish them clearly, incorrectly labeling them all as None. Additionally, aspects such as waiting space, waiting time, and waiting staff, where the boundary between service and atmosphere is ambiguous or falls into general anecdotes, are consistently mislabeled as None.
Furthermore, our threshold filtering that assigns a high-confidence Target None label for connotations with high similarity but lower errors supports a better distinction between semantic similarity and causal relationships. This is particularly true when excluding personalized terms and korean proper nouns like ‘Fixture (i.e., local native)’ used by reviewers. As a method to mitigate Spuriousness Haig (2003); Udomcharoenchaikit et al. (2022), we expect that our upper filtering will enable Target None predictions to function as data that helps the model learn the actual entailment relationships between sentences without relying solely on surface-level similarity in NLI pairs.
Thus, future research will delve into the association between embeddings from Bi-Encoder and Cross-Encoder models, and explore alternative methods for filtering or joint representation of translated NLI set. Simultaneously, the KR3 test set effectively capture real-world imbalances, deliberately selecting sentences with a broader vocabulary is necessary for enhanced diversity. Fully constructing the KR3 dataset through human annotation, it will serve as a benchmark for assessing LMs in Korean ABSA. Furthermore, Korean data from other domains should be acquired and evaluated.
Appendix F Computational Resources
All experiments in this study were conducted using high-performance computing resources from a leading cloud platform. The computational infrastructure included instances with 50GB or more of RAM, featuring NVIDIA A100 and V100 GPUs, chosen for advanced parallel processing capabilities essential for our LMs.