### v6

# KPC-cF: Aspect-Based Sentiment Analysis via Implicit-Feature Alignment with Corpus Filtering

Abstract
Investigations into Aspect-Based Sentiment Analysis (ABSA) for Korean industrial reviews are notably lacking in the existing literature. Our research proposes an intuitive and effective framework for ABSA in low-resource languages such as Korean. It optimizes prediction labels by integrating translated benchmark and unlabeled Korean data. Using a model fine-tuned on translated data, we pseudo-labeled the actual Korean NLI set. Subsequently, we applied LaBSE and MSP-based filtering to this pseudo-NLI set as implicit feature, enhancing Aspect Category Detection and Polarity determination through additional training. Incorporating dual filtering, this model bridged dataset gaps and facilitates feature alignment with minimal resources. By implementing alignment pipelines, our approach aims to leverage high-resource datasets to develop reliable predictive and refined models within corporate or individual communities in low-resource language countries. Compared to English ABSA, our framework showed an approximately 3% difference in F1 scores and accuracy. We will release our dataset and code for Korean ABSA, at this link [^1] https://anonymous.4open.science/r/KPC-cF-21E8.
Machine Learning, ICML
1 Introduction
In low-resource downstream tasks such as Korean ABSA, constraints exist in constructing ABSA systems that are socially and industrially beneficial (e.g., obtaining accurate labels and high-quality training data, building a efficient serving model). Addressing this challenge is fundamentally crucial for the practical implementation of multilingual ABSA leveraging the advantages of language models (Zhang et al., 2021; Lin et al., 2023). On the other hand, ABSA utilizing Large Language Models like ChatGPT can perform labeling through prompt tuning. However, due to reduced sensitivity to gradient-based tuning, the model may exhibit overconfidence in random labels, and limitations persist regarding resources for training and inference (Wang et al., 2023; Wu et al., 2023; Kossen et al., 2024).
(a) TR w/o PL-cF     (b) TR w/ PL-cF
Figure 1: t-SNE visualization of last [CLS] embeddings extracted from KR3 test set by two different experimental BaselineXLM-R encoders. Our $L_{\text{align}}$ on filtering data encourages the encoder to produce discriminable representations of different sentiment polarities, focusing on relevant aspects.
Figure 2: A diagram illustrating the two phase of our method: (1) Fine-tuning Kor-SemEval and generate pseudo-labeled KR3, (2) Fine-tuning KR3 using baseline model selected phase 1. We illustrated the filtering process (right) for fine-tuning KR3 data. Blue arrows (left & middle) indicate that this model is used to predict best label of review.
In this study, we derive pseudo-labels for real Korean reviews using machine-translated English ABSA data, inspired by the past research (Balahur & Turchi, 2012; Hoshino et al., 2024b). Moreover, we employ Dual filtering on the actual Korean corpus converted to implicit NLI task (Hendrycks & Gimpel, 2017; Sun et al., 2019; Feng et al., 2022), thereby constructing an effective framework coined as Korean ABSA using Pseudo-Classifier with Corpus Filtering (KPC-cF) to achieve implicit-feature alignment. Through this process, we assess the impact of our constructed classifier on the practical alignment of actual reviews. We validate that the pseudo-classifier, generated through the sentence-pair approach, outperforms the single approach in translation task. Furthermore, using the model that predicts the translated dataset most effectively as a baseline, we generate pseudo-labels for actual data and conduct real-world testing of Korean ABSA. This involves subsequent fine-tuning the filtered corpus based on language-agnostic embedding similarity for review and aspect sentence pairs, along with setting a threshold for Maximum Softmax Probability (MSP) in pseudo-labels.
The main contributions of our work are:
  • This is, to our knowledge, the first approach to generating a pseudo-classifier for automatic classification of aspect-based sentiment in the actual Korean domain.
  • For actual review-based ABSA, we propose a filtered NLI corpus as implicit feature and fine-tuning framework that enables important data selection in low-resource languages on models trained with high-resource dataset.
  • A new challenging dataset of Korean ABSA, along with a KR3 and translated benchmark correlated with cross-lingual understanding.
2 Task Description
In ABSA, Sun et al. (2019) set the task as equivalent to learning subtasks 3 (Aspect Category Detection) and 4 (Aspect Category Polarity) of SemEval-2014 Task 4 simultaneously. A similar approach was adopted for Korean ABSA in industrial reviews, aiming to develop a task-specific model through comparison of two PLMs (see Appendix§B), differing only in tokenization, vocabulary size, and model parameters (Conneau et al., 2020). Defining a unified-serving model using multi-label-multi-class classification from a task-oriented perspective was considered impractical due to challenges in modifying pre-training set and the ongoing injection of additional data, rendering implicit mapping unattainable (Sun et al., 2020; Ahmed et al., 2022; Qin & Joty, 2022; Kossen et al., 2024). Consequently, problem has been redefined into BERT-based tasks (see Appendix§B.1).
3 Two phase of Pseudo-Classifier
3.1 Motivation and Contribution
The goal of this research is to propose a framework for achieving the best ABSA on actual data with Korean nuances through high-resource languages. Past research by Balahur & Turchi (2012) has shown that Machine Translation (MT) systems can obtain training data for languages other than English in general sentiment classification. However, Li et al. (2020); Zhang et al. (2021) use explicitly aligned translation features, limiting the exploration of discriminative real target features in low-resource languages like Korean (Fei et al., 2022; Zhao et al., 2024). Also, although it was a different domain at Zhou et al. (2021), we found it necessary to investigate whether the concept of pseudo-labels could help bridge the gap of feature $\Phi$ between translated source data $\bm{{D_{S}}_{t}}$ and actual target language data $\bm{D_{T}}$. Therefore, we attempted the following two phases to assess the impact of the generated pseudo-classifier, fine-tuned using translated datasets from the ABSA benchmark and pseudo-labeled actual review data, on Korean ABSA. Figure 2 shows the two-phase pseudo-classifiers we will employ. In the first phase, we investigate the data structure exhibiting performance aligned with ${D_{S}}_{t}$ by thoroughly learning the fine patterns of the labeled benchmark for each model. NLI structure is referenced based on our research domain and language settings, following the prior work (Hoshino et al., 2024a). In Phase 2, we fine-tune the baseline $\Psi_{pre}(\Phi({D_{S}}_{t});\theta_{{D_{S}}_{t}})$, which was effective in training on ${D_{S}}_{t}$, by additionally incorporating pseudo-labeled $D_{T}$. Employing the tuned model $\Psi_{post}(\Phi(D_{T});\theta_{{{D_{S}}_{t}}\rightarrow D_{T}})$, we conduct predictions and evaluations on manually labeled actual Korean reviews. Throughout this process, Dual filtering is performed to enhance implicit features $\Phi(D_{T})$. Detailed description of our training strategy is provided in Section§3.4, Appendix§C.
Algorithm 1 Dual Filtering
1:  function Sampling($Target$, $\Psi_{pre}$)
2:   for $i=1$ to Target do
3:    $x_{s},x_{a}\leftarrow$ Target sample $i$       $\triangleright\,Attach\,aspect$
4:    $score_{L}\leftarrow$ LaBSE($x_{s},x_{a}$)
5:    Add $score_{L}$ to $temp$
[…]
7:    $score_{MSP}\leftarrow$ $\textit{MSP}_{\Psi_{pre}}$($y\mid x_{s},x_{a}$)
8:    if $score_{MSP}>threshold_{\tau_{1}}$
9:     Add $(x_{s},x_{a},y,score_{L})$ to $batch$
10:    end if
11:   end for
12:   $avg_{\tau_{2}}\leftarrow\frac{1}{N}\sum_{i=1}^{N}score_{L}[i]$ in $temp$
13:   $batch\leftarrow batch$ [$batch.score_{L}>avg_{\tau_{2}}$]
14:   return $\Phi(batch)$ for $\Psi_{post,\,joint}$ Tuning
15:  end function
3.2 LaBSE based Filtering
Our main approach, we aim to extract good-quality sentences-pair from the pseudo-NLI corpus. Language Agnostic BERT Sentence Embedding model (Feng et al., 2022) is a multilingual embedding model that supports 109 languages, including some Korean languages. Feng et al. (2022) suggested that the dual-encoder architecture of the LaBSE model, originally designed for machine translation in source-target language data (Batheja & Bhattacharyya, 2022, 2023), can be applied not only to other monolingual tasks like Semantic Textual Similarity (STS) but also to data (i.e., sentence pair-set) filtering for creating high-quality training corpora in terms of meaning equivalence. Therefore, to mitigate performance degradation caused by the linguistic gap between translated ${{D_{S}}_{t}}$ and actual Korean ${D_{T}}$ during fine-tuning, we introduce the following filtering method that enables the identification of meaning equivalence or connotation (Ghadery et al., 2019) in actual Korean sentence-pairs, even when viewed from the perspective of model trained on bilingual translation pairs. We generate the sentence embeddings for the review text and aspect of the pseudo-NLI corpora using the LaBSE model. Then, we compute the cosine similarity between the review text and aspect sentence embeddings. After that, we extract good quality NLI sentences based on a threshold value of the similarity scores.
LaBSE scoring
Let $D_{T}=\{(x_{s}^{i},x_{a}^{i})\}_{i=1}^{N}$ be a pseudo-NLI corpus with $N$ examples, where ${x_{s}^{i}}$ and ${x_{a}^{i}}$ represents $i^{th}$ review and aspect sentence respectively. We first feed all the review sentences present in the pseudo-parallel corpus as input to the LaBSE model [^2] https://huggingface.co/sentence-transformers/LaBSE, which is a Dual encoder model with BERT-based encoding modules to obtain review sentence embeddings ($S_{i}$). The sentence embeddings are extracted as the l2 normalized [CLS] token representations from the last transformer block. Then, we feed all the aspect sentences as input to the LaBSE model to obtain aspect sentence embeddings ($A_{i}$). We then compute cosine similarity $(score_{i})$ between the review and the corresponding aspect sentence embeddings.
$S_{i}=LaBSE\left(x_{s}^{i}\right)$ (1)
$A_{i}=LaBSE\left(x_{a}^{i}\right)$ (2)
$score_{i}=cosine\_similarity\left(S_{i},A_{i}\right)$ (3)
We aimed to apply the LaBSE scoring to the actual Korean dataset $D_{T}$, KR3, intending to facilitate flexible learning compared to the Eng-Kor translated dataset ${D_{S}}_{t}$, Kor-SemEval (see Section§4.1).
3.3 Confidence score Filtering
Meanwhile, we need to develop a classifier $\Psi_{post}$ capable of optimal predictions on the $D_{T}$ test, which can be considered as out-of-distribution data separate from the ${D_{S}}_{t}$. Drawing on previous research (Arora et al., 2021), we expect that language shifts (i.e., ${D_{S}}_{t}\leftrightarrow D_{T}$) embody both Background and Semantic shift characteristics. To ensure robust learning in both aspect detection and sentiment classification, we introduce additional thresholding on Maximum Softmax Probability (MSP; Hendrycks & Gimpel 2017) after LaBSE-based filtering on the $D_{T}$ train set. When considering an input $x=(x_{s},x_{a})\in\mathcal{X}$ and its corresponding pseudo-label $y\in\mathcal{Y}$, the score $s(x)$ for MSP is expressed as:
$s_{\textit{MSP}}(x)=\max_{k\in\mathcal{Y}}p_{\Psi_{\text{pre}}}(\hat{y}=k\mid x).$ (4)
Through this, we intended a dual scoring and filtering process to ensure that our $\Psi_{post}$ does not retrain on misplaced confidence or subpar prediction outcomes for out-of-distribution data. This process is defined by the following expected success rate of the trained models, where $\mathcal{F}_{\text{Dual}}$ filters the sentence pairs and pseudo-labels.
$P_{\Psi_{\text{post}}}=\mathbb{E}_{(x_{s},x_{a},\hat{y})\sim D_{T}}\left[\mathcal{F}_{\text{Dual}}(x_{s},x_{a},\hat{y})\right]$ (5)
Our algorithm for calculate scores and filter with the target $D_{T}$ batch set can be found in Algorithm 1.
3.4 Target-aligned Objective
In this section, we focus on the langauge adaptation of ABSA. The input data includes a set of labeled sentences from a translated source data ${D_{S}}_{t}=\{(x_{s}^{h},x_{a}^{h},y^{h})\}_{h=1}^{N}$ and a set of unlabeled sentences from a target language data $D_{T}=\{(x_{s}^{i},x_{a}^{i})\}_{i=1}^{N}$. Our goal is to validate our filtering technique across three types of tasks: (1) training only pseudo-labeled $D_{T}$, (2) joint training using both ${D_{S}}_{t}$ and $D_{T}$, (3) transfer learning from ${D_{S}}_{t}$ to $D_{T}$. Namely, unlike Chen et al. (2023), we maintain token embeddings and fine-tuned the transformer body & head. Contrary to Gururangan et al. (2020), we trained with ${D_{S}}_{t}$ and then reinitialized the classifier layer for further fine-tuning. This approach examines if the universal encoder and params $\theta$ aligned with ${D_{S}}_{t}$ enhance implicit alignment via pseudo-labels of refined $D_{T}$.
$\mathcal{L}_{pre}=-\frac{1}{N}\sum_{h=1}^{N}\mathbb{E}_{(x^{(h)},y^{(h)})\sim{D_{S}}_{t}}\left[\log p(y^{(h)}|x^{(h)};\theta)\right]$
$\mathcal{L}_{post}\!$ $=\!-\frac{1}{N}\sum_{i=1}^{N}\mathbb{E}_{p(x^{(i)},\hat{y}^{(i)})}\left[\log p(\hat{y}^{(i)}|x^{(i)};\theta^{\prime})\!\cdot\mathbb{I}(\hat{s}^{(i)}\geq\tau)\right]$
$\mathcal{L}_{align}=\mathcal{L}_{pre}+\mathcal{L}_{post}$ (6)
$\theta_{D_{T}}\leftarrow\theta^{\prime}-\hat{s}\cdot\nabla_{\theta^{\prime}}\mathcal{L}_{post}\quad\text{where}\;\;\theta^{\prime}=\theta_{body}+\theta_{cls}^{\text{init}}$ (7)
$\Psi_{\text{joint}}(\Phi({D_{S}}_{t}\sqcup D_{T});\theta_{{D_{S}}_{t}\sqcup D_{T}})$ was also evaluated, despite its inefficiency in terms of memory and computation, for data validation purposes. Thus, shuffling was used in $\Psi_{\text{joint}}$.
Assumption 3.1 (Contextual, Sementical gradient).
Reparameterization of Goyal et al. (2024) in Appendix§A , At timestep $t=i$, gradient of expected loss with respect to $W_{KQ,H}$ observes
$\theta_{C}^{\top}[-\nabla_{W_{KQ}}\mathcal{L}_{align}(W^{(i)})]\Phi(D_{T}\cdot\mathbb{I}(\hat{s}^{(i)}\geq\tau_{1,2}))>0,$
$\theta_{S}^{\top}[-\nabla_{W_{H}}\mathcal{L}_{align}(W^{(i)})]\Phi(D_{T}\cdot\mathbb{I}(\hat{s}^{(i)}\geq\tau_{1,2}))>0.$ (8)
Assumption 3.2 (Target-Aligned Feature).
Under Assumptions in objective, for any example $(x_{s},x_{a})\in$ $D_{T}$, after the gradient step at timestep $t=i$, the value embedding of the sentence-pair tokens is more predictive of the label $y$.
$\sigma\left(W_{H}^{(i)\top}W_{V}^{(i)}\Phi(x)\right)_{y}-\sigma\left(W_{H}^{(h)\top}W_{V}^{(h)}\Phi(x)\right)_{y}>0$ (9)
4 Experimental Setup
4.1 Dataset for Fine-Tuning and Test
Kor-SemEval
We translate the SemEval-2014 Task 4 (Pontiki et al., 2014) dataset [^3] http://alt.qcri.org/semeval2014/task4/. Moreover, it is evaluated for Korean ABSA. The training data was machine-translated (by Google Translate), and Test data was corrected manually only for fewer than 10 instances where abnormal translations occurred after machine translation. Each sentence contains a list of aspect ${x_{a}}$ with the sentiment polarity ${y}$. Ultimately, given a sentence ${x_{s}}$ in the sentence, we need to:
  • detects the mention of an aspect ${x_{a}}$;
  • determines the positive or negative sentiment polarity ${y}$ for the detected aspect.
This setting allows us to jointly evaluate Subtask 3 (Aspect Category Detection) and 4 (Aspect Category Polarity).
Model       #Training Task Type Aspect Category Polarity
Storage/Tokens Precision Recall Micro-F1 4-way acc 3-way acc Binary
GPT-4o mini (OpenAI, 2024)      -   - - 28.00 99.91* 43.74 84.41 86.13 89.10
Babel-9B-Chat (Zhao et al., 2025)      -   - - 64.31 87.88 74.27 75.66 76.57 79.37
Zero-Shot (Li et al., 2020) 1.43MB 0.38M $D_{S}$ 75.31 74.41 74.15 81.52 84.00 88.73
Trans-Ta (Zhang et al., 2021) 1.54MB 0.37M ${D_{S}}_{t}$ 94.04* 78.94 85.78 74.84 77.37 80.75
BaselinemBERT+PL 4.74MB 1.28M $D_{T}$ 89.71 77.60 83.21 77.96 80.15 84.01
BaselinemBERT+PL-cF 2.53MB 0.68M $D_{T}$ 87.45 76.28 81.46 78.49 80.69 84.70
BaselineXLM-R+PL 4.60MB 1.13M $D_{T}$ 90.58 79.26 84.54 83.36 85.71 89.88
BaselineXLM-R+PL-cF 2.15MB 0.52M $D_{T}$ 89.16 79.24 83.89 83.89 86.25 90.40
BaselinemBERT+TR+PL 6.28MB 1.73M ${D_{S}}_{t}\sqcup D_{T}$ 92.15 78.28 84.64 80.45 82.71 86.57
BaselinemBERT+TR+PL-cF 4.07MB 1.13M ${D_{S}}_{t}\sqcup D_{T}$ 90.15 81.20 85.42 80.88 83.15 86.96
BaselinemBERT+TR+PL 6.28MB 1.73M ${D_{S}}_{t}\rightarrow D_{T}$ 91.08 77.48 83.73 78.37 80.58 84.60
BaselinemBERT+TR+PL-cF 4.07MB 1.13M ${D_{S}}_{t}\rightarrow D_{T}$ 90.03 77.80 83.47 78.94 81.16 85.09
BaselineXLM-R+TR+PL 6.14MB 1.51M ${D_{S}}_{t}\sqcup D_{T}$ 91.88 83.66 87.57 83.43 85.78 89.34
KPC-cF 3.69MB 0.89M ${D_{S}}_{t}\sqcup D_{T}$ 92.72 83.30 87.75 83.52 85.90 89.56
BaselineXLM-R+TR+PL 6.14MB 1.51M ${D_{S}}_{t}\rightarrow D_{T}$ 91.50 84.23 87.69 84.21 86.57 90.35
KPC-cF 3.69MB 0.89M ${D_{S}}_{t}\rightarrow D_{T}$ 92.46 84.53 88.29 84.34 86.72 90.89
Table 1: KR3 test set results for Aspect Category Detection (middle) and Aspect Category Polarity (right). We reported the computational requirements as #Training. Specifically, for Baseline+PL, it refers to the overall capacity of structured NLI data and total number of valid tokens from KR3 Train. The performances where filtered models outperform non-filtered models under the same training conditions are highlighted in bold & underline. KPC-cF†, is jointly fine-tuned on the Kor-SemEval & Filtered KR3 Train. Task Type is represented by Section§3.4. For GPT-4o mini, Babel-9B-Chat, to reflect diversity in few-shot, we report the average of two results measured with temperature 0 or 0.7 and top-p 1, extending Wang et al. (2023). Zero-Shot, Trans-Ta (Li et al., 2020; Zhang et al., 2021) were evaluated using our settings in Appendix§C.
KR3
Unlike the domains previously used for Korean sentiment classification (Lee et al., 2020; Yang, 2021; Ban, 2022), Korean Restaurant Review with Ratings (KR3) is a comprehensive dataset encompassing various food service establishments, constructed from actual certified map reviews. In the case of restaurant reviews, words and expressions that evaluate positive and negative are mainly included, and real users often infer what a restaurant is like by looking at its reviews. Accordingly, Jung et al.3 constructed the KR3 dataset by crawling and preprocessing user reviews and star ratings of websites that collect broad food service information and ratings. KR3 has 388,111 positive and 70,910 negative, providing a total of 459,021 data plus 182,741 unclassified data, and distributed to Hugging Face [^4] https://huggingface.co/datasets/leey4n/KR3. We structured our training and test datasets to match the size of Kor-SemEval. Specifically, we addressed potential biases by randomly sampling indices from the original KR3, ensuring that evaluations for a specific restaurant were non-overlapping. Additionally, we maintained an even distribution of positive, negative, and neutral (ambiguous) classes, irrespective of the aspects indicated in the original KR3. This preprocessing step aimed to capture a comprehensive representation of sentiments across diverse attributes of sentences in the dataset. Subsequently, the data were configured to suit sentence pair classification. To allocate polarity labels for each aspect within the KR3 dataset, pseudo-labeling was conducted utilizing the optimal model identified during the Kor-SemEval performance evaluation. Pseudo-labels were assigned to the KR3 training data, and post pseudo-labeling, the test data underwent manual re-labeling by researchers.
Table 8 shows some Kor-SemEval and KR3 training data samples. In the case of KR3, the negative aspect is better reflected. Meanwhile, while Kor-SemEval gave neutrality to mediocre service, KR3 did not give neutrality to mediocre taste. While positive and negative data have been sufficiently accumulated and reflected, the tendency for a lack of neutral data can be confirmed in advance through some samples. We have organized both Kor-SemEval and KR3 data as open-source to facilitate their use in various training and evaluation scenarios.
5 Experiment
5.1 Main Experiment: KR3 Test Set
On the other hand, based on the results from Kor-SemEval, we examined the dissimilarity specific to the trans-align task between mBERT and XLM-RBase (see Appendix§D). Accordingly, We opted for the Baseline-NLI approach (i.e., BaselinemBERT, XLM-R), which demonstrated the best performance, as the base model for Phase 2 (see Figure 2, Section§3.1). We conducted evaluations on KR3 test data.
5.2 Main Results
To investigate the effect of features $\Phi(D_{T})$ for each corpus, we conduct baseline tuning comparisons between the PL and the PL-cF (see Table 1, 2 for details). The variants of our tuning framework includes:
(1) BaselinePLM+PL (Pseudo-Labeled data): Fine-tuning the Baseline PLM in an NLI manner with pseudo-KR3.
(2) BaselinePLM+PL-cF (Corpus Filtering): Fine-tuning Baseline PLM with the NLI data obtained by rejecting instance from pseudo-KR3, where the threshold of MSP (Hendrycks & Gimpel, 2017) is less than 0.5 and the cosine similarity between LaBSE embeddings is less than 0.15.
(3) BaselinePLM+TR (TRanslated data)+PL: Fine-tuning Baseline PLM in an NLI manner (pre-tuned or jointly fine-tuned with Kor-SemEval) using pseudo-KR3.
(4) KPC-cF (BaselinePLM+TR+PL-cF): Fine-tuning Baseline PLM in an NLI manner (pre-tuned or jointly fine-tuned with Kor-SemEval) using PL-cF.
Results on the KR3 test set are presented in Table 1 and Figure 5. We find that the KPC-cF approach achieved adequately trained results in both subtasks for the actual korean data. The model pre-tuned with Kor-SemEval achieves the best performance in Aspect Category Detection (ACD). For Aspect Category Polarity (ACP), it performs exceptionally well in the tuning of Pseudo-Labels, especially in the Binary setting. Filtered Pseudo-Labels preserve this characteristic well and amplify the performance of all metrics within ACP.
Dataset Count LaBSE. S Pred. Prob Corr.C
(mean) (mean) (target none)
[…]
Table 2: Number of instances, average scores, and standard deviations for 4 classes, as well as the correlation coefficient between two scores in the none class, for KR3 fine-tuning data. We set the model-specific dataset, PL and PL-cF.
Figure 3: Top - Maximum Probability Distribution of the Fine-Tuned Model, KPC-cF (left) vs. BaselineXLM-R+TR (right), Bottom - Maximum Probability Distribution of the BaselineXLM-R+TR with LaBSE Score Distribution, All classes (left) vs. 4 classes (right).
6 Discussion
In Phase 1, XLM-R, known for its proficiency in capturing cross-lingual representations, exhibits an underfitting tendency concerning the contextual disparities in aspect vocabulary within a single task. This can be attributed to data scarcity relative to model availability for each classifier or viewed as a limitation in single task using SPM in low-resource Korean ABSA (Son et al., 2023). Nevertheless, in the NLI task, it showcases potential by outperforming mBERT, guided by ”aspect”. Conversely, mBERT demonstrates stable results in both single and NLI tasks, exhibiting an overall accuracy increase, particularly in the NLI task (see Appendix§D). Furthermore, the pseudo-labels (PL) generated in Phase 2 refine polarity estimation, thereby enhancing classifier performance. Notably, PL-cF plays a crucial role in maintaining and improving accuracy and F1 score, even with fewer samples, unlike a simple addition of PL. The impact of further training on PL (PL further trained on TR) is evident through embedding analysis and ablation studies based on thresholding (see Figure 1, 5). Additionally, GPT-4o mini exhibits high recall but significantly lower precision, as it assigns polarity-labeled aspects to a much larger set. In contrast, Trans-Ta achieves higher precision, suggesting that alignment models trained on translated data can refine contextual categories relevant to domain classification. This further implies the importance of conservative evaluation in review classification, particularly for aspects that may impact reputation, as well as its significance in real-time inference.
Model Task Type KorSTS KLUE-STS
Zero-shot Inference
BaselineKLUE-R pre-trained 23.78 -10.91
BaselineKLUE-R-Large pre-trained 13.33 -18.17
SetFitKLUE-R-13domain pre & fine-tuned 29.49 8.37
[…]
BaselineXLM-R+TR ${D_{S}}_{t}$ 10.26 2.55 *
BaselineXLM-R+ PL $D_{T}$ 12.18 7.59
BaselineXLM-R+TR+PL ${D_{S}}_{t}\rightarrow D_{T}$ 9.21 6.35
KPC-cF ${D_{S}}_{t}\rightarrow D_{T}$ 16.26 20.50
Table 3: Contextual Implication Comparisons of intermediate layer features as mean-pooled embeddings (Choi et al., 2021) on general STS. We report Spearman’s $\rho$ (%). Higher is better. *Values are for p $\approx$ 0.1; others within p $<$ 0.05. SetFit is selected based on the number of fine-tuning data. [^6] For details on SetFitKLUE-R, refer to Tunstall et al. (2022) and https://huggingface.co/mini1013.
Looking at Figure 3, compared to BaselineXLM-R+TR, KPC-cF significantly increases the maximum probability of Target Polarity through additional fine-tuning with filtered data. Despite a 40% data reduction, it prevents biased class learning and maintains a balanced probability distribution across Target classes. Furthermore, for sentiment polarity data where the target is detected, higher LaBSE scores correspond to increased model confidence. KPC-cF also exhibits scalability in intermediate and shallow-layer feature comparisons. It enhances the model’s ability to capture general contextual relationships between sentence embeddings (see Table 3), as confirmed through zero-shot embedding evaluations using correlation coefficients against Ground Truth. Additionally, training a linear model with word-level features further validates the influence of filtered data characteristics on training outcomes (see Table 4). This confirms effectiveness of our approach in refining semantic label information. Our framework improves model stability and reliability by optimally rejecting uncertain or biased data.
Model Feature Span TR+PL $\rightarrow$ $\Delta$ TR+PL-cF
${D_{S_{t}\sqcup T}}$ ${x_{s\oplus a}}$ Micro-F1 4-way acc
XLM-RBase I.T + LR S.T +16.01 +6.29
mBERT I.T + LR S.T +13.14 +5.19
TF-IDF I.T + LR - +0.19 +0.19
Table 4: Sementic Stability Performance variation of TR+PL vs. TR+PL-cF (generated by BaselineXLM-R) utilizing Logistic Regression as a shallow feature classifier for different Input Token IDs. We combined each Data & NLI format w/w.o special tokens. To assess the contribution of previous ${D_{S}}_{t}$, shuffling was omitted.
6.1 Further Analysis
To analyze the types of anticipatory effects that enable language models to exhibit improved performance in filtering corpora, we conducted a qualitative analysis by examining 116 cases where KPC-cF outputs matched Ground Truth but not those of the misaligned model (i.e., BaselineXLM-R+TR+PL). Interestingly, we identified the following four patterns that complement the prediction of misaligned model: (1) False Detection Refining: misaligned model incorrectly predicted aspects not mentioned in the reviews with a specific polarity, but KPC-cF refined these to ‘None’. (2) Mismatch Refining: misaligned model incorrectly predicted aspects mentioned in the reviews with a completely different polarity, which KPC-cF refined to the correct polarity. (3) Multi Context Detecting: KPC-cF detected contexts within a review with multiple aspects, correctly assigning the ‘None’ overlooked by misaligned model to the correct polarity. (4) Single Context Detecting: KPC-cF detected contexts within a single-aspect review and correctly assigned the polarity that misaligned model had overlooked as ‘None’.
Patterns 1 and 2 emphasize the role of KPC-cF in providing error feedback. The filtered PL corpus indicates the presence of more significant polarities, guiding responses from the model. In contrast, patterns 3 and 4 demonstrate that the filtering framework allows the language model to utilize aligned aspect information. This adaptation of cohesive contexts from the translation corpus optimizes problem-solving for the target task. The frequency of each pattern from our qualitative assessment is provided in Table 5. On the other hand, given the somewhat limited effectiveness compared to Refining, the Detecting patterns suggest that training initiated from TR still possesses both scalability limits and potential. Therefore, we propose additional analysis and directions in Section§7.
Observed Instances %
(gt-matched, not $\hat{y}_{\text{TR+PL}}$)
False Detection Refining 34.38%
Mismatch Refining 31.25%
Multi  Context Detecting 18.75%
Single Context Detecting 15.62%
Table 5: Qualitative assessment results of KPC-cF predictions on the KR3 test set; We analyze aligned model outcomes with Refining and Detecting patterns.
Figure 4: Left & Middle - Maximum Probability Distribution of the BaselinemBERT+TR with LaBSE Score Distribution, All classes (left) vs. 4 classes (middle), Right - Maximum Probability Distribution of the Fine-Tuned Model, BaselinemBERT+TR.
7 Limitations and Future Work
Acknowledging limitations is crucial before delving into future work. While mBERT exhibits somewhat inadequate transfer effects, which can be attributed to its smaller model size. Due to pre-training solely on translation data, the scores are densely populated, resulting in limited performance improvement through additional training. Nevertheless, unlike jointly with Target None, the regression line of maximum probability for Target Polarity still trends upward according to LaBSE score (see Table 2, Figure 4), indicating that ongoing evaluation of threshold points for the two scores can be done amidst more diverse and precise patterns. Meanwhile, we conducted direct inference on individual NLI inputs through a demo [^7] https://huggingface.co/KorABSA. Upon detecting various aspects within the reviews, we observed a slight predominance towards the ”None” class within the class-wise similar probability distributions per sentence in TR training, while a more concentrated prediction tendency towards the binary class was observed in PL training. Furthermore, a tendency for one class to dominate over two or more aspect classes within review sentences was identified. This appears to stem from the class imbalance in TR and contextual differences (Jin et al., 2023) within all review data (i.e., TR & PL).
Our analysis of the PL-cF thresholding reveals that while it reduces the number of predicted Target None classes resulting from extensive training on None class in TR, it also qualitatively explains why the error rate for Target None can be lower (see Table 7, 7). Contrary to experimental results that show it is reasonable to predict high-confidence Target Polarity labels when there is high surface similarity between sentences, cases where sentence similarity is less than 0.15 but predicted as Target None with confidence greater than 0.5 can hinder the model’s understanding of deeper semantic relationships. For instance, issues related to hygiene and cleanliness are not detected through the overall atmosphere, and in cases where multiple aspects overlap, the model fails to distinguish them clearly, incorrectly labeling them all as None. Additionally, aspects such as waiting space, waiting time, and waiting staff, where the boundary between service and atmosphere is ambiguous or falls into general anecdotes, are consistently mislabeled as None. Furthermore, our threshold filtering assigns a high-confidence Target None label for connotations with high similarity but lower errors, which supports a better distinction between semantic similarity and causal relationships. This is particularly effective when excluding personalized terms and Korean proper nouns, such as ‘Fixture (i.e., local native)’ used by reviewers. Our current translation-based training does not sufficiently capture the tone of sentences composed of proper nouns and particles from various domains within Korea. However, we have confirmed through various evaluations (Table 1, 2, 3, 4, 5, 7, 11 and Figure 5, 6, 7, 8) that our filtering approach aids in learning actual level of entailment relationships, moving beyond merely identifying similar relationships between sentences. To mitigate spurious correlations (Haig, 2003; Udomcharoenchaikit et al., 2022), we propose upper filtering that enables Target predictions to function as data facilitating the model’s learning of true entailment relationships between sentences, without relying solely on surface-level similarity in NLI pairs.
[…]
Thus, future research will delve into the association between embeddings from Bi and Cross-Encoder models, and explore alternative methods for filtering or joint representation of translated NLI set. Simultaneously, the KR3 test set effectively capture real-world imbalances, deliberately selecting sentences with a broader vocabulary is necessary for enhanced diversity. Fully constructing the KR3, it will serve as a benchmark for assessing LMs in Korean ABSA. Furthermore, Korean data from other domains should be acquired and evaluated. From an industrial perspective, our model could be integrated with multilingual AI and agents (Bi et al., 2024; OpenAI, 2025) as a refining module.
8 Conclusion
Aspect-Based Sentiment Analysis (ABSA) is a key subfield in text analytics, but the lack of high-quality labeled data hinders its industrial development. This study aims to address the language gap in ABSA. By aligning implicit features through dual filtering, we propose an framework to align target languages solely through translation of existing benchmarks. This approach is expected to optimize tasks for both Korean and low-resource languages with limited pre-training on rich native data. Additionally, we presented Kor-SemEval, KR3 train (pseudo labeled & filtered), and test data (Gold) composed of fine-grained set. We invite the community to extend Korean ABSA by providing new datasets, trained models, evaluation results, and metrics.
Acknowledgements
We thank Chanseo Nam for manually correcting and annotating the Kor-SemEval and KR3 testset used to test our model, during his undergraduate research. We also thank Kyeongpil Kang, Byungok Han, Woohan Yun, Soyoung Yang, Joosun Yum and Guijin Son for their valuable discussions. Thanks to DAVIAN, MODULABS and UST AI as well. This research was supported by Brian Impact Foundation, a non-profit organization dedicated to the advancement of science and technology for all.
Appendix A Formulation under Assumption: Enhanced NLI-Based Pseudo-Classifier
Similarity and Confidence
The similarity between $X_{s}$ (review tokens) and $X_{a}$ (aspect tokens) is defined probabilistically:
$\text{sim}(X_{s},X_{a})=\mathbb{E}_{p(X_{s},X_{a})}\left[\frac{X_{s}\cdot X_{a}}{\|X_{s}\|\|X_{a}\|}\right],$ (10)
where $p(X_{s},X_{a})$ represents the joint distribution of the input pairs. The confidence $C$ of target polarity is derived as:
$C=f\big{(}\text{sim}(X_{s},X_{a})\big{)},$ (11)
where $f$ is a monotonically increasing function, such as $f(x)=\log(1+x)$ or $f(x)=\sigma(ax+b)$.
Relationship with Sementic Label Accuracy
The accuracy of pseudo label $Y$, denoted as $Acc_{Y}$, is modeled as an expectation (Arora et al., 2021) over the confidence:
$Acc_{Y}=\mathbb{E}_{p(C)}[g(C)],$ (12)
where $g(C)$ is a monotonically increasing function, e.g., $g(C)=kC$ for $k>0$.
Loss Function and Gradients
The model’s loss $\mathcal{L}(\theta)$ is expressed as:
$\mathcal{L}(\theta)=\mathbb{E}_{p(X_{s},X_{a},Y)}\big{[}\ell(Y,\hat{Y})\cdot f(\text{sim}(X_{s},X_{a}))\big{]},$ (13)
where $\ell(Y,\hat{Y})$ denotes the loss between the predicted label $\hat{Y}$ and the pseudo label $Y$. The gradients of the loss function are given by:
$\nabla_{\theta}\mathcal{L}(\theta)=\mathbb{E}_{p(X_{s},X_{a},Y)}\big{[}\nabla_{\theta}\ell(Y,\hat{Y})\cdot f(\text{sim}(X_{s},X_{a}))\big{]}.$ (14)
Updates to Model Parameters
The output layer weights $W_{H}$, responsible for sementic label-specific information, are updated as:
$\Delta W_{H}=-\eta\nabla_{W_{H}}\mathcal{L}(W),$ (15)
where $-\nabla_{W_{H}}\mathcal{L}(W)$ represents the gradient direction, and $\eta$ is the learning rate. Similarly, the Key ($W_{KQ}$) weights are updated as:
$\Delta W_{KQ}=-\eta\nabla_{W_{KQ}}\mathcal{L}(W).$ (16)
Gradient Propagation to Tokens
The gradient propagation from the loss $\mathcal{L}(\theta)$ affects the token embeddings $X_{s}$ and $X_{a}$, refining their representations during training:
$\nabla_{X_{s}}=\frac{\partial\mathcal{L}(\theta)}{\partial X_{s}},\quad\nabla_{X_{a}}=\frac{\partial\mathcal{L}(\theta)}{\partial X_{a}}.$ (17)
This ensures that the context similarity $\text{sim}(X_{s},X_{a})$ evolves dynamically, aligning with the label $Y$.
Alignment and Overall Dynamics
The alignment of gradients and confidence in the context of label-specific updates can be expressed as:
$\theta_{Context}^{\top}\big{[}-\nabla_{W_{KQ}}\mathcal{L}_{align}(W)\big{]}\Phi(D_{T}\cdot\mathbb{I}(\hat{s}\geq\tau_{1,2}))>0,$ (18)
$\theta_{Sementic}^{\top}\big{[}-\nabla_{W_{H}}\mathcal{L}_{align}(W)\big{]}\Phi(D_{T}\cdot\mathbb{I}(\hat{s}\geq\tau_{1,2}))>0,$ (19)
where $\mathbb{I}(\hat{s}\geq\tau_{1,2})$ is an indicator function, $\Phi$ denotes a non-linear mapping, and $D_{T}$ represents task-specific adjustments.
This formulation highlights the relationship between token similarity ($X_{s},X_{a}$), parameter updates ($W_{H},W_{KQ}$), and their alignment with confidence and task-specific refinements, providing insights into the dynamic learning process.
Extension of Experimental Results
Figure 5: Performance of ACD and ACP on KR3 test set during adaptive fine-tuning ($D_{T}$ and ${D_{S}}_{t}\rightarrow D_{T}$). Left: results with the addition of other fine-tuned BaselineXLM-R. th denotes the threshold for confidence of pseudo-labeling, and L denotes the threshold for LaBSE filtering; Right: BaselineXLM-R tuning compared in this paper. Blue line represents KPC-cF.
A.1 Gradient Stability and Directionality from Dual Scoring
Notation. Let $x_{s},x_{a}\in\mathbb{R}^{d}$ denote the sentence embeddings in a compact metric space $\mathcal{E}$. Let $\Phi(x_{s},x_{a})\in\mathcal{E}\subset\mathbb{R}^{d}$ be the fused input representation. Let $\theta=\{W_{H},W_{V},W_{KQ}\}$ be the trainable parameters. Define:
$\text{sim}(x_{s},x_{a})$ $=\frac{\langle x_{s},x_{a}\rangle}{\|x_{s}\|\cdot\|x_{a}\|}\,\in[-1,1],$
$C(x_{s},x_{a})$ $=\max_{k}\,p(k\mid x_{s},x_{a};\theta)\,\in[0,1],$
$\mathcal{D}_{f}$ =$=\left\{(x_{s},x_{a},\hat{y})\in\mathcal{D}\,\middle|\,\text{sim}(x_{s},x_{a})>\tau_{1},\;C(x_{s},x_{a})>0.5\right\}$
where $\tau_{1}=\mathbb{E}_{\mathcal{D}}[\text{sim}]$ is the empirical mean similarity. The fixed confidence threshold of $0.5$ reflects the theoretical minimum for non-random predictions and ensures robustness under calibration skewness.
Let the empirical loss be:
$\mathcal{L}(\theta)=\frac{1}{|\mathcal{D}_{f}|}\sum_{(x_{s},x_{a},\hat{y})\in\mathcal{D}_{f}}\ell(f_{\theta}(x_{s},x_{a}),\hat{y})$
where $\ell$ is the cross-entropy loss.
Assumptions. (1) $\mathcal{E}$ is compact. (2) sim and $C$ are Lipschitz-continuous. (3) $\ell$ is locally convex and differentiable.
Lemma 1. Gradient Stability via Score Concentration
Let $L$ be the Lipschitz constant of sim.
$|\text{sim}(x_{1},x_{2})-\text{sim}(x_{1},x_{3})|\leq L\cdot\|x_{2}-x_{3}\|_{2}$
After filtering, we observe:
$\mu_{\text{sim}}^{f}$ $:=\mathbb{E}_{\mathcal{D}_{f}}[\text{sim}]>\mathbb{E}_{\mathcal{D}}[\text{sim}]=:\mu_{\text{sim}},$
$\sigma_{\text{sim}}^{f}$ $:=\sqrt{\mathbb{E}_{\mathcal{D}_{f}}[(\text{sim}-\mu_{\text{sim}}^{f})^{2}]}<\sqrt{\mathbb{E}_{\mathcal{D}}[(\text{sim}-\mu_{\text{sim}})^{2}]}=:\sigma_{\text{sim}},$
$\mu_{C}^{f}$ $:=\mathbb{E}_{\mathcal{D}_{f}}[C]>\mathbb{E}_{\mathcal{D}}[C]=:\mu_{C},$
$\sigma_{C}^{f}$ $:=\sqrt{\mathbb{E}_{\mathcal{D}_{f}}[(C-\mu_{C}^{f})^{2}]}<\sqrt{\mathbb{E}_{\mathcal{D}}[(C-\mu_{C})^{2}]}=:\sigma_{C}$
Define filtered gradient:
$\nabla_{\theta}\mathcal{L}_{f}=\frac{1}{|\mathcal{D}_{f}|}\sum_{(x_{s},x_{a},\hat{y})\in\mathcal{D}_{f}}\nabla_{\theta}\ell(f_{\theta}(x_{s},x_{a}),\hat{y})$
Score concentration implies bounded gradient deviation:
$\forall(x_{s},x_{a})\in\mathcal{D}_{f},\quad\mathbb{E}\left[\|\nabla_{\theta}\ell(f_{\theta}(x_{s},x_{a}),\hat{y})-\nabla_{\theta}\mathcal{L}_{f}\|_{2}^{2}\right]\leq\varepsilon_{1}(\sigma_{\text{sim}}^{f},\sigma_{C}^{f})$
Lemma 2. Confidence-Guided Gradient Direction under Reinitialization
Let the classifier weights be initialized as $W_{H}^{(t_{0})}\sim\mathcal{N}(0,\sigma^{2})$, and let $(x_{s},x_{a},y)\in\mathcal{D}_{f}$ be a filtered sample with high confidence $C(x_{s},x_{a})>\tau$. Then, at the training step $t=t_{0}$, where the classification head is reinitialized, the gradient satisfies:
$\left\langle\nabla_{W_{H}}\ell(f_{t_{0}}(x_{s},x_{a}),y),\Phi(x_{s},x_{a})\right\rangle>0$
indicating that gradient updates are initially aligned with the semantic direction $\Phi(x_{s},x_{a})$. This alignment facilitates faster convergence during early fine-tuning.
Lemma 3. Compactness and Gradient Directionality
Let $\{\Phi(x_{s}^{t},x_{a}^{t})\}_{t=1}^{\infty}\subset\mathcal{E}$ be a sequence of fused representations, where $\mathcal{E}$ is the compact representation space defined in Assumption (1).
By the topological definition of compactness, for every open cover $\{U_{\alpha}\}_{\alpha\in A}$ of $\mathcal{E}$, there exists a finite subcover. Hence, the sequence $\{\Phi(x_{s}^{t},x_{a}^{t})\}$ cannot diverge or oscillate unboundedly, and must lie within a finite union of open sets. This guarantees the existence of a convergent subsequence:
$\Phi(x_{s}^{t_{k}},x_{a}^{t_{k}})\rightarrow\Phi^{*}\in\mathcal{E}\quad\text{as}\quad k\rightarrow\infty.$
Assuming $\ell$ is continuous in $\Phi$ and $f_{\theta}$ is differentiable, we obtain:
$\nabla_{\theta}\ell(f_{\theta}(\Phi(x_{s}^{t_{k}},x_{a}^{t_{k}})),y_{t_{k}})\rightarrow\nabla_{\theta}\ell(f_{\theta}(\Phi^{*}),y^{*}).$
Furthermore, let $q^{(t)}:=W_{Q}\Phi(x_{s}^{t},x_{a}^{t})$ and $k^{(t)}:=W_{K}\Phi(x_{s}^{t},x_{a}^{t})$ denote the query and key projections. By linearity and the convergence $\Phi(x_{s}^{t},x_{a}^{t})\to\Phi^{*}$, we have:
$\langle q^{(t)},k^{(t)}\rangle=\langle W_{Q}\Phi(x_{s}^{t},x_{a}^{t}),W_{K}\Phi(x_{s}^{t},x_{a}^{t})\rangle\rightarrow\langle W_{Q}\Phi^{*},W_{K}\Phi^{*}\rangle.$
Therefore, the attention score matrix defined by
$A^{(t)}:=\text{softmax}\left(\frac{q^{(t)}(k^{(t)})^{\top}}{\sqrt{d_{k}}}\right)$
converges to
$A^{*}:=\text{softmax}\left(\frac{W_{Q}\Phi^{*}(W_{K}\Phi^{*})^{\top}}{\sqrt{d_{k}}}\right),$
and the attention structure becomes stable across samples.
Finally, the gradient w.r.t. $W_{KQ}$ also converges:
$\nabla_{W_{KQ}}\ell(f_{\theta}(\Phi(x_{s}^{t},x_{a}^{t})),y_{t})\rightarrow\nabla_{W_{KQ}}\ell(f_{\theta}(\Phi^{*}),y^{*}),$
with directional alignment preserved:
$\left\langle\nabla_{W_{KQ}}\mathcal{L}_{f},\;W_{KQ}\Phi^{*}\right\rangle>0.$
Lemma 4. Logit Monotonicity Post-Filtering
Let $\sigma$ denote softmax and define logit scores:
$P_{\text{pre}}(y|x_{s},x_{a})$ $=\sigma(W_{H}^{(t_{0})\top}W_{V}^{(t_{0})}\Phi(x_{s},x_{a}))_{y},$
$P_{\text{post}}(y|x_{s},x_{a})$ $=\sigma(W_{H}^{(t)\top}W_{V}^{(t)}\Phi(x_{s},x_{a}))_{y}$
Then:
$P_{\text{post}}(y|x_{s},x_{a})-P_{\text{pre}}(y|x_{s},x_{a})>\delta>0$
indicating alignment with Assumption 3.2.
Theorem. Convergence under Low Gradient Dispersion
Assume:
$\forall(x_{s},x_{a})\in\mathcal{D}_{f},\quad\mathbb{E}\left[\|\nabla_{\theta}\mathcal{L}-\mathbb{E}[\nabla_{\theta}\mathcal{L}]\|^{2}\right]\leq\sigma^{2}$
Then for $L$-smooth $\mathcal{L}$ and Adam optimizer:
$\lim_{t\to\infty}\left\|\frac{1}{|\mathcal{D}_{f}|}\sum_{(x_{s},x_{a},\hat{y})}\nabla_{\theta}\ell(f_{\theta}(x_{s},x_{a}),\hat{y})\right\|_{2}\to 0$
Summary of Theoretical Guarantees
The dual-filtered set $\mathcal{D}_{f}$ satisfies:
[…]
Under the NTK regime (Jacot et al., 2018), these enhancements extend to shallow-layer generalization (Table 4), improving decision boundary sharpness.
Figure 6: Multi-Head Attention Visualization. In intermediate layers associated with domain and linguistic understanding (Clark et al., 2019), the proposed KPC-cF model exhibits stronger self-attention activations across heads for tokens relevant to target language relations. In the final encoder layer responsible for classification, KPC-cF shows greater focus on label-relevant information in the [CLS] embedding, with reduced influence from irrelevant noise. This visualization is consistent with Lemma Lemma 3. Compactness and Gradient Directionality.
Appendix B Background and Related Work
Real-World ABSA
Early research in ABSA primarily focused on single-element prediction, such as aspect term, category extraction (Yin et al., 2016; Xu et al., 2018; Ma et al., 2019; Bu et al., 2021), and sentiment polarity classification (Wang et al., 2016a; Huang & Carley, 2018; Hu et al., 2019; Zhang & Qian, 2020). After the advent of pre-trained language models, BERT-based architectures have become dominant in ABSA research (Zhang et al., 2022; Cai et al., 2023; Gao et al., 2022; Wang et al., 2024). Additionally, text-to-text or LLM approaches (Gou et al., 2023; Yang et al., 2024) have been explored to improve ABSA tasks. These approaches leveraging pre-trained models assume high-performance extraction in settings with sufficient monolingual context and labeled data. However, in real-world multilingual ABSA scenarios, challenges such as paraphrasing variations, semantic consistency with aspects, and polarity prediction quality (Zhang et al., 2021; Hsu et al., 2021; Fei et al., 2022) can arise, leading to performance degradation. To address these challenges in real-world scenarios, we employ an NLI-based pseudo-labeling and filtering approach. By inferring sentiment through structured and refined sentence relationships, our method mitigates semantic shifts and polarity distortions while adapting to the target data. This approach enhances (1) robust alignment between review-aspect context and sentiment information, even when word alignment is disrupted during translation, and (2) generalization performance in multilingual settings.
mBERT
Multilingual BERT is a BERT trained for multilingual tasks. It was trained on monolingual Wikipedia articles in 104 different languages. It is intended to enable mBERT finetuned in one language to make predictions for another. Azhar & Khodra (2020) and Jafarian et al. (2021) show that mbert performs effectively in a variety of multilingual Aspect-based sentiment analysis. It is also actively used as a base model in other tasks of Korean NLP (Lee et al., 2021; Park et al., 2021), but is rarely confirmed in Korean ABSA tasks. Thus, our study used the pre-trained mBERT base model with 12 layers and 12 heads (i.e., 12 transfomer encoders). This model generates a 768-dimensional vector for each word. We used the 768-dimensional vector of the Extract layer to represent the comment. Like the English language subtasks, a single Dense layer was used as the classification model.
XLM-R
XLM-RoBERTa (Conneau et al., 2020) is a cross-lingual model that aims to tackle the curse-of-multilingualism problem of cross-lingual models. It is inspired by RoBERTa (Liu et al., 2019), trained in up to 100 languages, and outperforms mBERT in multiple cross-lingual ABSA benchmarks (Zhang et al., 2021; Phan et al., 2021; Szołomicka & Kocon, 2022). However, like mBERT, Korean ABSA has yet to be actively evaluated, so we used it as a base model. We use the base version (XLM-RBase) coupled with an attention head classifier, the same optimizer as mBERT.
B.1 Classification approach
Single sentence Classification
BERT for single-sentence classification tasks. For ABSA, we fine-tune the pre-trained BERT model to train $n_{a}$ (i.e., number of aspect categories) classifiers for all aspects and then summarize the results. The input representation of the BERT can represent a pair of text sentences in a sequence of tokens. In the case of a single task, only one review text is tokenized and inputted as a sequence of tokens. A given token’s input representation is constructed by summing the corresponding token, segment, and position embeddings. For classification tasks, the first word of each sequence is a unique classification embedding [CLS]. Segment embeddings in single sentence classification use one.
Sentence-pair Classification
Based on the auxiliary sentence constructed as aspect word text, we use the sentence-pair classification approach to solve ABSA. The input representation is typically the same with the single-sentence approach. The difference is that we have to add two separator tokens [SEP], the first placed between the last token of the first sentence and the first token of the second sentence. The other is placed at the end of the second sentence after its last token. This process uses both segment embeddings (In the case of XLM-RoBERTa, an additional one is placed in the first position, resulting in a total of three segment embeddings). For the training phase in the sentence-pair classification approach, we only need to train one classifier to perform both aspect categorization and sentiment classification. Add one classification layer to the Transformer output and apply the softmax activation function.
Ensemble
Meanwhile, we additionally use a voting-based ensemble, a typical ensemble method. The ensemble can confirm generalized performance based on similarity of model results in NLI task (Xu et al., 2020). So, we add separate power-mean ensemble result to identify a metric that amplifies probabilities based on the Pre-trained Language Models (PLMs). we reported the ensemble results of the top-performing models trained on NLI tasks for each PLM.
Kor-SemEval Train Aspect Polarity
서비스는 평범했고 에어컨이 없어서 가격 (price) 없음 (None)
편안한 식사를 할 수 없었습니다. 일화 (anecdotes) 없음 (None)
음식 (food) 없음 (None)
(The service was mediocre and the lack of air conditioning 분위기 (ambience) 부정 (Negative)
made for a less than comfortable meal.) 서비스 (service) 중립 (Neutral)
KR3 Train Aspect Polarity
가로수길에서 조금 멀어요 점심시간에 대기 엄청납니다 가격 (price) 없음 (None)
일행 모두 있어야 들어갈 수 있어요 맛은 보통이에요 일화 (anecdotes) 부정 (Negative)
음식 (food) 없음 (None)
(It’s a little far from Garosu-gil. There’s a huge wait during lunch time. 분위기 (ambience) 부정 (Negative)
You have to have everyone in your group to get in. The taste is average.) 서비스 (service) 없음 (None)
Table 8: Samples of Kor-SemEval and KR3 train dataset.
Appendix C Experimental Setup
C.1 Hyperparameter
All experiments are conducted on two pre-trained cross-lingual models. The XLM-RoBERTa-base and BERT-base Multilingual-Cased model are fine-tuned. The number of Transformer blocks is 12, the hidden layer size is 768, the number of self-attention heads is 12, and the total number of parameters for the XLM-RoBERTa-base model is 270M, and BERT base Multilingual-Cased is 110M. When fine-tuning, we keep the dropout probability at 0.1 and the initial learning rate is 2e-5. In the Preliminary Experiment (Appendix§D), we set the batch size to 3, this setting was applied to both single and NLI tasks. The max length was set to 512, and for epochs beyond 3, no significant performance improvement was observed, so the results from epoch 2 were noted. Subsequently, in Main Experiment (Section§5.2), we fine-tuned with a batch size of 16, following the pattern of previous experiments (Karimi et al., 2021). All models trained on Korean data achieved faster convergence and optimal performance. As no significant improvements were observed beyond 10 epochs, results were reported at 4 epochs. Each reported metric is the average of runs with four different random seeds to mitigate the effects of random variation in the results.
C.2 Metrics
The benchmarks for SemEval-2014 Task 4 are the several best performing systems in Sun et al. (2019), Wang et al. (2016b) and Pontiki et al. (2014). When evaluating Kor-SemEval and KR3 test data with subtask 3 and 4, following Sun et al. (2019), we also use Micro-F1 and accuracy respectively.
Appendix D Preliminary Experiment
D.1 Experiment: Kor-SemEval
We conducted evaluations for each of the mBERT-single, XLM-RBase-single, mBERT-NLI, XLM-RBase-NLI, and NLI-ensemble models. As there is no officially converged dataset and model research specifically for Korean ABSA, We included the results from the previous SemEval14 research and Kor-SemEval to compare and evaluate the performance in Korean.
Model SemEval-14
Precision Recall Micro-F1
BERT-single 92.78 89.07 90.89
BERT-pair-NLI-M 93.15 90.24 91.67
Models trained & evaluated on Kor-SemEval
mBERT-single 92.16 77.95 84.46
XLM-RBase-single 91.01 49.37 64.01
mBERT-NLI 91.10 79.90 85.14
XLM-RBase-NLI 91.37 83.71 87.37
NLI-ensemble 93.70 81.27 87.04
Table 9: Test set results for Aspect Category Detection. We use the results reported in BERT-single and BERT-pair-NLI-M (Sun et al., 2019) for English dataset together with our results.
Model SemEval-14
4-way acc 3-way acc Binary
BERT-single 83.7 86.9 93.3
BERT-pair-NLI-M 85.1 88.7 94.4
Models trained & evaluated on Kor-SemEval
mBERT-single 68.20 71.84 79.52
XLM-RBase-single 62.93 66.29 75.20
mBERT-NLI 73.95 77.90 84.87
XLM-RBase-NLI 79.41 83.66 89.98
NLI-ensemble 78.24 82.43 89.65
Table 10: Test set accuracy (%) for Aspect Category Polarity. We use the results reported in BERT-single and BERT-pair-NLI-M (Sun et al., 2019) for English dataset together with our results.
D.2 Preliminary Results
Results on Kor-SemEval are presented in Table 10 and Table 10. Similar to the SemEval results, it was confirmed that tasks converted to NLI tasks tend to be better than single tasks, with mBERT achieving better results in single and XLM-RBase in NLI. The XLM-RBase-NLI model performs best, excluding precision for aspect category detection. It also works best for aspect category polarity. The NLI-ensemble model was the best in precision but performed poorly in other metrics.
Appendix E Attention Head Visualization
For token visualization (Figure 6, 7, 8), we used input sentences from the KR3 dataset, structured as [CLS] review [SEP] aspect [SEP]. An example snippet in Figure 6 is as follows: review = “맛 집이라고 찾아서 갔는데 손님이 오든 말든 불친절에 물컵에 기름기가 잘잘 흐르고 위생도 꽝이고 면에서는 밀가루 냄새만 나고 먹다가 다 남기고 실망하고 나왔어요. 왜 이런 집이 맛 집으로 평가받는지 모르겠다.”, aspect = “음식”.
Appendix F Computational Resources
All experiments presented in this study were conducted using high-performance computing resources provided by a leading cloud platform. The computational infrastructure utilized for the experiments included instances with 50GB or more of RAM and GPUs, specifically the NVIDIA A100 and V100 models. The choice of these GPUs was made to harness the advanced parallel processing capabilities essential for the computational demands of our LMs.
Model GPU.t Thruput PL Filtering PL Alignment Aspect Category Polarity
LaBSEavg MSP0.5 METEOR SemScore Precision Recall Micro-F1 4-way acc 3-way acc Binary
BaselineXLM-R+TR+PL 41 30.45K 48.93 65.85 91.5 84.23 87.69 84.21 86.57 90.35
BaselineXLM-R+TR+PL 31 22.00K 48.84 66.62 91.63 85.51 88.44 83.31 85.66 89.78
BaselineXLM-R+TR+PL 38 29.53K 48.99 65.75 90.39 84.18 87.14 84.50 86.88 90.86
KPC-cF 29 21.30K 49.03 66.58 92.46 84.53 88.29 84.34 86.72 90.89
Table 11: Ablation Study LaBSE-based similarity thresholding improves aspect detection, while confidence-based filtering via MSP enhances polarity classification. The proposed KPC-cF achieves complementary effects by combining both filters, yielding robust performance with reduced data. Throughput denotes the total number of training samples processed across both pre- and second-stage. The joint-trained KPC-cF † required 47 minutes of GPU time. Evaluation of PL sample quality using METEOR and SemScore (Aynetdinov & Akbik, 2024). Dual-filtered samples in KPC-cF show better input-label alignment than unfiltered PL data.
Figure 7: Multi-Head Attention Visualization. In intermediate layers associated with domain and linguistic understanding (Clark et al., 2019), the proposed KPC-cF model exhibits stronger self-attention activations across heads for tokens relevant to target language relations. In the final encoder layer responsible for classification, KPC-cF shows greater focus on label-relevant information in the [CLS] embedding, with reduced influence from irrelevant noise. This visualization is consistent with Lemma Lemma 3. Compactness and Gradient Directionality. Snippet is as follows: review = “맛 집이라고 찾아서 갔는데 손님이 오든 말든 불친절에 물컵에 기름기가 잘잘 흐르고 위생도 꽝이고 면에서는 밀가루 냄새만 나고 먹다가 다 남기고 실망하고 나왔어요. 왜 이런 집이 맛 집으로 평가받는지 모르겠다.”, aspect = “서비스”.
Figure 8: Multi-Head Attention Visualization. In intermediate layers associated with domain and linguistic understanding (Clark et al., 2019), the proposed KPC-cF model exhibits stronger self-attention activations across heads for tokens relevant to target language relations. In the final encoder layer responsible for classification, KPC-cF shows greater focus on label-relevant information in the [CLS] embedding, with reduced influence from irrelevant noise. This visualization is consistent with Lemma Lemma 3. Compactness and Gradient Directionality. Snippet is as follows: review = “편안한 느낌의 토스트 킴. 식사 한 끼로 좋은 토스트를 팝니다. 고기를 좋아해서 갈 때마다 시키는 건 불고기, 삼겹살, 돈가스, 소고기 치즈 토스트 등… 각자의 매력이 있고 한국식 토스트인데 정말 두둑합니다… 단품 세트로 시키면 감자튀김에 음료도 나오니 웬만한 패스트푸드 보다 낫다 보시면 되겠네요… 오천 원 하나로 두둑이 먹을 수 있어요.”, aspect = “가격”.